OpenStudy (anonymous):

how do i factor quadratics when the leading coefficiant isnt 1x EX: 9X-6x+1

OpenStudy (amistre64):

a few ways .... quadratic equation might be the simplest, and by simplest I mean not the simplest lol

OpenStudy (anonymous):

wats a way other than quadratic formula

OpenStudy (amistre64):

other ways are to set up the factors, if they can be integered like this: (x+m/a) (x+n/a) such that: ax^2 +bx + c = 0 factors as: a*c = m*n m+n = b

OpenStudy (amistre64):

reduce m/a and n/a as far as they will go; and if they are still a fraction, put the bottom in front. (ax + m) (ax + n)

OpenStudy (amistre64):

if we use your example we have: 9x^2-6x+1 = 0 9(1) = 9, factors of n are: 1,9; 9+1 = 10 -3,-3; -3-3 = -6 (x-3/9) (x-3/9), simplify (x-1/3) (x-1/3) ; put the bottoms in front (3x-1) (3x-1), or simply (3x-1)^2

OpenStudy (amistre64):

*factors of 9 are ... :)

OpenStudy (anonymous):

thanks takes me awhile but ill get it